Chris starts his run at 7:00 a.m. He runs at a rate of 4 miles per hour. His course takes him three miles out and then back along the same path. John runs at a rate of 6 miles per hour, starting at 7:10 a.m., and he runs the same course. How many minutes elapse between the time John finishes and the time Chris finishes?
Answer by texttutoring(324)

We will use the equation d=vt, where d=distance, v=velocity (or speed), and t=time.
Chris:
d = 3 + 3 = 6 miles
v = 4 mph
t = ?
d=vt
6=4*t
t=6/4
t=1.5 hours
John:
d = 3 + 3 = 6 miles
v = 6 mph
t = ?
d=vt
6=6*t
t=6/6
t=1 hour
Chris started at 7:00am, so he will end at 8:30am.
John started at 7:10am, so he will end at 8:10am.
20 minutes will elapse after John finishes.